OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5). A 52 B -1.5 C 22 D 102

OpenStudy (hoblos):

f(5) = 2(5)^2 +2 = 2(25) +2 = 52

OpenStudy (hoblos):

A